The area of a rectangular field is 153sqm.If its length is 8m longer than its breadth,find its width.
Answer by LinnW(1048)

Set L = length , W = width
L = W + 8
area = L*W or
area = (W + 8)(W)
153 = (W + 8)(W)
153 = W^2 + 8W
add -153 to each side
0 = W^2 + 8W -153
Notice that 153 = 3*51 = 3*3*17 or 9*17
0 = (W + 17)(W - 9)
So W = -17 or W = 9
W needs to be positive so the width is 9m
and since L = W + 8, length = 9 + 8 = 17m